# When does bottom-up beat top-down in hierarchical community detection?

### v2

@@ 6 Numerical Results
In this section, we investigate the empirical performance of different HCD strategies on synthetic and real networks. More precisely, we compare the bottom-up approach of Algorithm 1 with other HCD algorithms. To implement Algorithm 1, we use spectral clustering with the Bethe-Hessian matrix, as proposed in DCT (21), as the initial clustering step. This is a spectral method that does not require the knowledge of the number of communities $K$ and that was shown to work very well on synthetic and real data sets [^8] The code for this algorithm is available online at https://lorenzodallamico.github.io/codes/.. In all the following, bottom-up refers to Algorithm 1, top-down to the recursive bi-partitioning algorithm of LLB+ (22). To facilitate reproducible research, our code is available on GitHub [^9] https://github.com/dakrod/bottom_up_HCD.
@@ 6.1 Synthetic Data Sets
We compare bottom-up and top-down approaches on synthetic data sets by computing the accuracy at different depths. We define accuracy at depth $q$ as $1-\mathrm{loss}\left(\mathrm{sc}(q,\mathcal{C},\mathcal{T}),\mathrm{sc}(q,\widehat{\mathcal{C}},\widehat{\mathcal{T}})\right)/N$, where $\mathrm{loss}$ is given by (3.1) and $\mathrm{sc}(q,\mathcal{C},\mathcal{T})$ (resp., $\mathrm{sc}(q,\widehat{\mathcal{C}},\widehat{\mathcal{T}})$) are the ground truth (resp., predicted) super-communities defined in (4.3).
@@ 6.1.1 Binary Tree SBMs
We first generate synthetic BTSBMs with depth $d$, where the probability of an edge between two nodes whose lowest common ancestor has depth $k$ is $p_{k}=a_{k}\log N/N$. We compare the accuracy of top-down and bottom-up at each depth. We show in Figure 2 the results obtained on BTSBMs with depth $3$, with $200$ nodes in each primitive community (thus $N=2^{3}\times 400=3200$ nodes in total). We let $a_{0}=40$, $a_{3}=100$, and the values of $a_{1}$ and $a_{2}$ vary in the range ($a_{0},a_{3}$), with the condition $a_{1}<a_{2}$. Solid lines in each panel show the exact recovery threshold of the given method at different depths. We observe the strong alignment between the theoretical guarantees and the results of the numerical simulations.
Figure 2: Performance of bottom-up and top-down algorithms on BTSBMs with depth 3, $N=3200$ nodes, and interaction probabilities $p_{k}=a_{k}\log N/N$ where $a_{0}=40$ and $a_{3}=100$, as a function of $a_{1}$ and $a_{2}$. We vary $a_{1}\leq a_{2}$ from $a_{0}$ to $a_{3}$. The empirical performance of the algorithms is measured by the accuracy at each depth, given by the color scale (results are averaged over 10 realizations). Large circles represent exact recovery (i.e., perfect accuracy on each of the 10 runs), and small crosses represent a non-exact recovery. The colored solid lines delimit the theoretical exact recovery thresholds for each algorithm on the various depths (given by Equations (5.4)-(5.5)); for a given depth $q$, these equations provide a single condition for bottom-up, but $q$ conditions for top-down. At depths 1 and 2, the regimes where exact recovery can be achieved are the areas above the solid line(s). At depth 3, the area lies below the threshold drawn by the red line (and above the blue and green lines for top-down; this area forms a small triangle).
6.1.2 Ternary Tree SBMs
We first show in Figure 3 the dendrograms and trees obtained by top-down and bottom-up algorithms. Although both algorithms generate binary trees, ternary structures appear in Figures 3(b) and 3(c). Nevertheless, we observe in Figure 3(c) that the dendrograms obtained by the top-down algorithm show some inversions.
(a) Ground truth tree
(b) bottom-up
(c) top-down
@@ 6.1.3 Robustness of Linkage to Misclustering Errors
We numerically test the robustness of the average-linkage procedure, when the expected number of mistakes done in the clustering of the primitive communities is $\eta N$, with $\eta>0$. Consider the two following scenarios described in Section 3.3:
  • a uniform scenario in which $\widetilde{\mathcal{C}}_{\mathrm{uniform}}$ is sampled so that the misclustered nodes are assigned to a cluster uniformly at random among all clusters (this is the setting of Example 1);
  • an adversarial scenario in which $\widetilde{\mathcal{C}}_{\mathrm{adversarial}}$ is sampled so that the misclustered nodes are assigned to a cluster chosen uniformly at random among the clusters that are the furthest away from their true cluster (this is the setting of Proposition 1).
We denote by $\widehat{\mathcal{T}}_{\mathrm{uniform}}$ (resp., by $\widehat{\mathcal{T}}_{\mathrm{adversarial}}$) the tree obtained by the average-linkage algorithm from the bottom communities $\widetilde{\mathcal{C}}_{\mathrm{uniform}}$ (resp., from $\widetilde{\mathcal{C}}_{\mathrm{adversarial}}$). To assess the correct recovery of the tree, we denote by $S\left(\mathcal{T},\mathcal{C}\right)$ the $N$-by-$N$ tree similarity matrix. The elements of this matrix are defined, for $i\in C_{a}$ and $j\in C_{b}$ (with $a,b\in\mathcal{L}_{\mathcal{T}}$), by
$\Big{(}S\left(\mathcal{T},\mathcal{C}\right)\Big{)}_{ij}\ =\ |\mathrm{lca}_{\mathcal{T}}(a,b)|,$ (6.1)
where $|\mathrm{lca}_{\mathcal{T}}(a,b)|$ is the depth of the lowest common ancestor of $C_{a}$ and $C_{b}$ in the tree $\mathcal{T}$. The error in recovering the tree is then defined as
@@
$r_{s}\ =\ \frac{\left\|S\left(\widehat{\mathcal{T}}_{s},\widetilde{\mathcal{C}}_{s}\right)-S\left(\mathcal{T},\widetilde{\mathcal{C}}_{s}\right)\right\|_{F}^{2}}{\left\|S\left(\mathcal{T},\widetilde{\mathcal{C}}_{s}\right)\right\|_{F}^{2}},$
where $s\in\{\text{uniform},\text{adversarial}\}$. This ratio equals 00, as long as the algorithm succeeds in restoring the tree completely, even if the bottom communities $\widetilde{\mathcal{C}}_{s}$ contain mistakes. We plot in Figure 5 the values of these ratios. As established in Example 1, we see that the tree is correctly recovered in the uniform scenario. For the adversarial scenario, we observe the failure of the tree recovery when the proportion of misclustered nodes $\eta$ becomes too large. Nonetheless, when $\eta$ is small, the average-linkage procedure still recovers the tree correctly. This highlights the robustness of the average-linkage and is in alignment with Proposition 1.
(a) Uniform mistakes
(b) Adversarial mistakes
Figure 5: Recovery from error done by the average-linkage procedure on a a balanced BTSBM with depth 3, $N=4000$, and $p_{k}=0.08\beta^{3-k}$, as a function of $\beta$ and of the misclustering error rate $\eta$. To distinguish exact recovery ($r_{s}=0$) from almost exact recovery, square symbols are used for the former, and circles for the latter. Results are averaged over 25 realizations.
@@ 6.2.1 High-School Contact Dataset
We illustrate the performance of HCD algorithms on a data set of face-to-face contacts between high school students. The data set was collected from the high school Lycée Thiers in Marseilles, France MFB (15) and is available at http://www.sociopatterns.org/. The 9 communities are classes of 327 students, and the weighted interactions are the number of proximity encounters during 5 school days. We can expect this network to possess hierarchical relationships between communities. Indeed, although each student belongs to one of the nine classes, these classes also define the student’s specialty. More precisely, there are four specializations:“MP” stands for mathematics and physics, “PC” for physics and chemistry, “PSI” for engineering, and “BIO” for biology. Furthermore, there are three “MP” classes (MP, MP*1, and MP*2), two “PC” classes (PC and PC*), one “PSI” class (PSI*), and three “BIO” classes (2BIO1, 2BIO2, and 2BIO3). Therefore, we expect that the hierarchy reveals the students’ specialties at a higher depth and the students’ classes at a lower depth.
The results of the two HCD algorithms are shown in Figure 6. The bottom-up algorithm always detects $31$ communities. In contrast, out of 100 runs, the top-down algorithm predicts on average 8.93 communities (7 runs detect $8$ communities and the remaining 93 runs detect 9 communities). The predictions of both algorithms are aligned with the ground truth. Indeed, the adjusted mutual information (AMI) VEB (09) between the nine classes taken as ground truth labels and the predicted communities is $0.938$ for bottom-up [^10] We compute the AMI with respect to the configuration obtained after merging the 31 bottom communities into 9 super-communities. and $0.945$ for top-down. We observe that bottom-up successfully recovers the class and specialization structure but, in addition, finds smaller communities. This likely represents small groups of friends inside a class: the bottom-up algorithm reveals a community structure that is richer than the ground truth itself.
Figure 6: Bottom-up and top-down algorithms on the high school data set. Vertices correspond to the students, the colors to the true classes, and edges of the graph are in grey. The hierarchical tree is drawn in black, and its root is marked by a star symbol.
@@ 6.2.2 Power Grid
We next consider the power grid of continental Europe from the map of the Union for the Coordination of Transmission of Electricity (UCTE) [^11] http://www.ucte.org. We use the same data set that was previously used for HCD SDYB (12). Figures 7 and 8 show the outputs of the bottom-up and top-down, respectively. We observe a higher correlation to geographical positions in the output of bottom-up. Moreover, the dendrogram obtained by top-down shows a significant amount of inversions, contrary to the one obtained by bottom-up.
@@ 7 Conclusion
In this paper, we have quantified this fundamental advantage within a class of random graph models (HSBM). We have proven that the latent tree of an HSBM can be recovered under weaker conditions than in the literature (average degree scaling as $\omega(1)$). Moreover, we have established that super-communities of intermediate depths could be exactly recovered up to the information-theoretic threshold, thus improving the previously known conditions for top-down algorithms. Finally, we have shown that the theoretic advantage of bottom-up carries over to relevant scales and real-world data. Both on synthetic and on real data sets, bottom-up HCD achieves better performance than top-down HCD.
@@
Appendix A Proofs of Section 3
@@ A.1 Proof of Theorem 1
With the assumption of this theorem, $\widehat{\mathcal{C}}$ is an almost exact estimator of the true clusters $\mathcal{C}$. Let $a\neq b\in\mathcal{L}_{\mathcal{T}}$, and recall the definition of the edge density $\rho(\cdot,\cdot)$ in (2.1). Lemma 2 implies that with high probability the edge density between the estimated clusters $\widehat{C}_{a}$ and $\widehat{C}_{b}$ is $\rho(\widehat{C}_{a},\widehat{C}_{b})=(1+o(1))p_{ab}$. Since the network is assortative and the hierarchy is non-flat, this implies that the first step of the linkage finds the two closest clusters, say $\widehat{C}_{a_{1}}$ and $\widehat{C}_{a_{2}}$. Then we see that for $b\in\mathcal{L}_{\mathcal{T}}\backslash\{a_{1},a_{2}\}$ we have $\mathrm{lca}(a_{1},b)=\mathrm{lca}(a_{2},b)$ and hence $p_{a_{1}b}=p_{a_{2}b}$. Therefore, from (2.2),
@@
Let us bound the two terms $E_{1}$ and $E_{2}$ on the right-hand side of (A.1) separately. To handle the first term, we will use [17, Lemma 4.1]. Let $\bar{p}=2N^{-1}(N-1)^{-1}\sum_{i<j}\operatorname{Var}(A_{ij})$. Because $\operatorname{Var}(A_{ij})=\Theta(\delta_{N})$, we have $\bar{p}=\Theta(\delta_{N})$. Moreover, because $\delta_{N}=\omega(N^{-1})$, we have $\bar{p}\geq 9N^{-1}$ for $N$ large enough. Therefore [17, Lemma 4.1] ensures that with probability at least $1-e^{3}5^{-N}$ we have
[…]
Applying Grothendieck’s inequality [17, Theorem 3.1], we obtain
[…]
where $c$ is Grothendieck’s constant, verifying $20<c<$ 2. Using (A.2) with $X_{i}=\hat{Z}_{ia}$ and $Y_{j}=\hat{Z}_{jb}$ leads to
@@
$\ \leq\ \left(\left|\widehat{C}_{b}\right|+\left|\widehat{C}_{a}\right|\right)\mathrm{loss}\left(\mathcal{C},\widehat{\mathcal{C}}\right),$
@@
A.2 Proof of Section 3.3
In this section, we prove the Theorem 2 and Proposition 1. We first start in Section A.2.1 by stating a lemma that gives a general condition for the robustness of average-linkage with respect to errors present in $\widetilde{\mathcal{C}}$. We then show in Sections A.2.2 and A.2.3 how to prove Theorem 2 and Proposition 1 from this Lemma 3. The main ingredients of the proof of Lemma 3 are given in Section A.2.4, whereas the most tedious computations are detailed in Section A.2.5.
@@ A.2.1 A general lemma
$B(h)$ =$\ =\ \begin{cases}2^{d-h-1}&h\in\{0,1,\cdots,d-1\},\\ 1&h=d.\end{cases}$ (A.4)
@@ A.2.2 Proof of Theorem 2
Because of assortativity, $p_{h_{2}}-p_{h_{1}}>0$ for all $h_{2}>h_{1}$ and $\bar{p}_{h}>p_{h_{ac}}$ for all $h>h_{ac}$. Moreover, for all $h_{1}<h_{2}$ such that $\zeta(h_{1})\ \leq\ \zeta(h_{2})$, we have $\zeta(h)-\zeta(h_{ac})\geq 0$. Therefore, for all $h_{ac}\geq d-2$, we have
@@
[…]
@@
A.2.3 Proof of Proposition 1

### v3

@@ 6 Numerical Results
In this section, we investigate the empirical performance of different HCD strategies on synthetic and real networks. More precisely, we compare Algorithm 1 (bottom-up) to the recursive bi-partitioning algorithm of li2022hierarchical (top-down). To obtain the flat clustering input for Algorithm 1, we use spectral clustering with the Bethe-Hessian matrix, as proposed in dall2021unified. This spectral algorithm does not require prior knowledge of the number of communities $K$ and has been shown to perform well on both synthetic and real datasets. [^7] The code for this algorithm is available online at https://lorenzodallamico.github.io/codes/. Moreover, recent theoretical work has established that the Bethe-Hessian matrix can be used to recover both the number of communities and the communities (stephan2024community). Our code is available on GitHub at https://github.com/daichikuroda/bottom_up_HCD.
@@ 6.1 Synthetic Data Sets
[…]
@@ 6.1.1 Binary Tree SBMs
We first generate synthetic BTSBMs of depth $d$, where the probability of an edge between two nodes whose lowest common ancestor has depth $k$ is $p_{k}=a_{k}\log N/N$. We compare the accuracy of top-down and bottom-up at each depth. We show in Figure 2 the results obtained on BTSBMs of depth $3$, with $400$ nodes in each primitive community (thus $N=2^{3}\times 400=3200$ nodes in total). We let $a_{0}=40$, $a_{3}=100$, and the values of $a_{1}$ and $a_{2}$ vary in the range ($a_{0},a_{3}$), with the condition $a_{1}<a_{2}$. Solid lines in each panel show the exact recovery threshold of the given method at different depths. We observe the strong alignment between the theoretical guarantees and the results of the numerical simulations.
Figure 2: Performance of bottom-up and top-down algorithms on BTSBMs of depth 3, $N=3200$ nodes, and interaction probabilities $p_{k}=a_{k}\log N/N$, where $a_{0}=40$ and $a_{3}=100$, as a function of $a_{1}$ and $a_{2}$. We vary $a_{1}\leq a_{2}$ from $a_{0}$ to $a_{3}$. The empirical performance of the algorithms is measured by the accuracy at each depth, given by the color scale (results are averaged over 10 realizations). Large circles represent exact recovery (i.e., perfect accuracy on each of the 10 runs), and small crosses represent a non-exact recovery. The colored solid lines delimit the theoretical exact recovery thresholds for each algorithm on the various depths (given by Equations (5.4)-(5.5)); for a given depth $q$, these equations provide a single condition for bottom-up, but $q$ conditions for top-down. At depths 1 and 2, the regimes where exact recovery can be achieved are the areas above the solid line(s). At depth 3, the area lies below the threshold drawn by the red line (and above the blue and green lines for top-down; this area forms a small triangle).
@@ 6.1.2 Robustness of Linkage to Misclustering Errors
Finally, we evaluate the performance of Algorithm 1 when the number of misclassified nodes increases. We generate BTSBMs with full and balanced trees of depth $3$ with 200 nodes in each community ($N=1600$ in total), and interaction probabilities given by $p_{k}=0.08\beta^{3-k}$, where $\beta$ is a parameter varying between 0 and $1$. At the two extremes, $\beta=0$ leads to a graph with $K$ communities disconnected from each other, while $\beta=1$ leads to an Erdős-Rényi graph without community structure.
Figures 3(a) and 3(b) show two confusion matrices, which are $K$-by-$K$ matrices where each element $(i,j)$ is $|C_{i}\cap\widehat{C}_{\tau^{*}(j)}|$, for two different average degree. We observe that mislabeled nodes are typically assigned to a community close to their true one. For instance, a misclassified node from community 0 may be incorrectly clustered into community 1 or 2, but it is highly unlikely to be placed in community 7. To reinforce this observation and further evaluate the validity of Assumption 3, we estimate the probability of misclassification at level $\ell\in[K]$, corresponding to the quantity $\zeta(\ell)$ in Assumption 3, by
@@
$\hat{\zeta}(\ell)\ =\ \frac{1}{N}\sum_{k=1}^{K}\sum_{m=k}^{K}\mathbb{1}\{|\mathrm{lca}(C_{k},C_{m})|=\ell\}\cdot\left|C_{k}\cap\widehat{C}_{\tau^{*}(m)}\right|,$
where $\tau^{*}=\operatorname*{arg\,min}_{\tau\in\mathcal{S}_{[K]}}\sum_{k=1}^{K}\left|C_{k}\,\Delta\,\widehat{C}_{\tau(k)}\right|$ is the optimal permutation of cluster labels. Figure 3(c) shows that, for any given $\beta$, we have $\hat{\zeta}(0)<\hat{\zeta}(1)<\hat{\zeta}(2)<\hat{\zeta}(3)$, showing that $\hat{\zeta}$ is non-decreasing with $\ell$. This observation aligns with Assumption 3. Finally, Figure 3(d) shows the tree recovery success rates with and without graph-splitting. A successful tree recovery occurs when the following condition is satisfied:
$\prod_{k=1}^{K}\prod_{m=k+1}^{K}\mathbb{1}\left\{|\mathrm{lca}(C_{k},C_{m})|\ =\ |\mathrm{lca}(\mathcal{C}_{\tau^{*}(k)},\mathcal{C}_{\tau^{*}(m)})|\right\}.$
The legend indicates the value of $\gamma$ used for graph splitting, which represents the proportion of edges allocated to estimating bottom community labels via the flat clustering algorithm, while the remaining edges are used to construct the hierarchy with Algorithm 1. We observe that Algorithm 1 maintains a high tree recovery rate even when the estimated bottom clusters contain many misclustered nodes. When the expected degree is $5$, Figure 3(c) shows that $\hat{\zeta}(d)=\hat{\zeta}(3)\approx 0.6$, meaning that approximately 40% of the nodes are misclustered. Nevertheless, Figure 3(d) indicates that Algorithm 1 achieves a tree recovery success rate of around 50%. This demonstrates the robustness of Algorithm 1 to a high proportion of misclustered nodes. Furthermore, while we introduced graph splitting for the theoretical analysis, our empirical results show that it may not be needed in practice.
(a) Average degree 55
(b) Average degree 10
(c) Types of errors
(d) Tree recovery rate
Figure 3: BTSBMs with depth 3, $N=1600$, and $\beta=0.3$. Figures 3(a) and 3(b) show two confusion matrices when the expected degree equals 5 and 10. Figure 3(c) shows the evolution of $\zeta$ as a function of the expected degree.Figure 3(d) shows the tree recovery success rate with and without graph splitting. Results of Figures 3(c) and 3(d) are averaged over 200 realizations.
@@ 6.2.1 High-School Contact Dataset
We evaluate HCD algorithms on a dataset of face-to-face interactions among 327 high school students from Lycée Thiers in Marseilles, France Mastrandrea_Fournet_Barrat_2015 (http://www.sociopatterns.org/). The network consists of 9 class-based communities, with weighted edges representing proximity encounters over five days. These classes also correspond to four academic specializations: mathematics & physics (MP), physics & chemistry (PC), engineering (PSI), and biology (BIO). The MP and BIO groups contain three sub-classes each, while PC has two and PSI one. The hierarchical structure should reflect specialties at higher depths and individual classes at lower depths.
The results of both HCD algorithms are shown in Figure 4. The bottom-up algorithm consistently detects 31 communities, while the top-down algorithm predicts an average of 8.93 (8 communities in 7 runs, 9 in 93 runs). Both align well with the ground truth, with adjusted mutual information (AMI) scores of 0.938 for bottom-up [^8] AMI is computed after merging the 31 bottom-up communities into 9 super-communities. and 0.945 for top-down. Notably, bottom-up recovers both class and specialization structures while also detecting smaller subgroups, likely representing groups of friend inside a same class, revealing a richer hierarchical structure than the ground truth itself.
Figure 4: Bottom-up and top-down algorithms on the high school data set. Nodes correspond to the students, colors to the true classes, and edges of the graph are in grey. The hierarchical tree is drawn in black, and its root is marked by a star symbol.
@@ 6.2.2 Power Grid
We next consider the power grid of continental Europe from the map of the Union for the Coordination of Transmission of Electricity (UCTE). We use the same data set that was previously used for HCD schaub2012markov. Figures 5 and 6 show the outputs of the bottom-up and top-down, respectively. We observe a higher correlation to geographical positions in the output of bottom-up. Moreover, the dendrogram obtained by top-down shows a significant amount of inversions, contrary to the one obtained by bottom-up.
@@ 7 Conclusion
In this paper, we have quantified this fundamental advantage within a class of random graph models (HSBM). We have proven that the latent tree of an HSBM can be recovered under weaker conditions than in the literature (average degree scaling as $\omega(1)$). Moreover, we have established that super-communities of intermediate depths could be exactly recovered up to the information-theoretic threshold, thus improving upon the previously known conditions for top-down algorithms. Finally, we have shown that the theoretic advantage of bottom-up carries over to relevant scales and real-world data. Both on synthetic and on real data sets, bottom-up HCD achieves better performance than top-down HCD.
Acknowledgements
This work was supported in part by the Swiss National Science Foundation (SNSF) under grant IZBRZ2_186313.
@@ Acknowledgements
Appendix A Proofs for Section 3
@@ A.1 Proof of Theorem 1
[…]
@@
Let us bound the two terms $E_{1}$ and $E_{2}$ on the right-hand side of (A.1) separately. To handle the first term, we will use [guedon2016community, Lemma 4.1]. Let $\bar{p}=2N^{-1}(N-1)^{-1}\sum_{i<j}\operatorname{Var}(A_{ij})$. Because $\operatorname{Var}(A_{ij})=\Theta(\delta_{N})$, we have $\bar{p}=\Theta(\delta_{N})$. Moreover, because $\delta_{N}=\omega(N^{-1})$, we have $\bar{p}\geq 9N^{-1}$ for $N$ large enough. Therefore [guedon2016community, Lemma 4.1] ensures that with probability at least $1-e^{3}5^{-N}$ we have
$\sup_{s,t\in\{-1,1\}^{n}}\left|\sum_{i,j}\left(A_{ij}-\mathbb{E}A_{ij}\right)s_{i}t_{j}\right|\ \leq\ 3N\sqrt{N\bar{p}}.$
Applying Grothendieck’s inequality [guedon2016community, Theorem 3.1], we obtain
$\sup_{\begin{subarray}{c}X_{1},\cdots,X_{N}\\ Y_{1},\cdots,Y_{N}\\ \forall i\in[N]\colon\|X_{i}\|_{2}\leq 1\\ \forall j\in[N]\colon\|Y_{j}\|_{2}\leq 1\end{subarray}}\left|\sum_{i,j}\left(A_{ij}-\mathbb{E}A_{ij}\right)X_{i}^{T}Y_{j}\right|\ \leq\ 3cN\sqrt{N\bar{p}},$ (A.2)
where $c$ is Grothendieck’s constant, verifying $0<c<2$. Using (A.2) with $X_{i}=\hat{Z}_{ia}$ and $Y_{j}=\hat{Z}_{jb}$ leads to
@@
[…]
@@
A.2 Proofs for Section 3.3
In this section, we prove the Theorem 2 and Proposition 1. We first start in Section A.2.1 by stating a lemma that gives a general condition for the robustness of average-linkage with respect to errors present in an estimator $\widetilde{\mathcal{C}}$ of $\mathcal{C}$, where we assume that $\widetilde{\mathcal{C}}$ is independent of the graph $G$. We then show in Sections A.2.2 and A.2.3 how to prove Theorem 2 and Proposition 1 from this Lemma 3. The main ingredients of the proof of Lemma 3 are given in Section A.2.4, whereas the most tedious computations are detailed in Section A.2.5.
@@ A.2.1 A General Lemma
$B(h)$ $\ =\ \begin{cases}2^{d-h-1}&h\in\{0,1,\cdots,d-1\},\\ 1&h=d.\end{cases}$ (A.4)
@@ A.2.2 Proof of Theorem 2
Because of assortativity, $p_{h_{2}}-p_{h_{1}}>0$ for all $h_{2}>h_{1}$ and $\bar{p}_{h}>p_{h_{ac}}$ for all $h>h_{ac}$. Moreover, we have $\zeta(h_{1})\ \leq\ \zeta(h_{2})$ and $\zeta(h)-\zeta(h_{ac})\geq 0$ for all $h_{1}<h_{2}$ and $h>h_{ac}$. Therefore, for all $h_{ac}\leq d-2$, we have
@@
$+$ $\sum_{h_{1}=h_{ac}+1}^{d-1}B^{2}(h_{1})\left(\zeta(h_{1})-\zeta(h_{ac})\right)^{2}\left(\bar{p}_{h_{1}+1}-p_{h_{ac}}\right)+\left(p_{d-1}-p_{h_{ac}}\right)\left(\zeta(d)-\zeta(h_{ac})\right)^{2}\ \geq\ 0,$
@@
A.2.3 An Adversarial Setting where $\zeta$ is not Non-decreasing
$\zeta_{\mathrm{adversarial}}(h)\ =\ \begin{cases}\frac{\eta}{2^{d-1}}&\text{ if }h=0,\\ 0&\text{ if }h\in\{1,\cdots,d-1\}.\end{cases}$ (A.8)
Proposition 1.
[…]